If the demand during review cycle is equal to 12, the demand during lead time is 11, and the safety stock is equal to 8, what is the order point?

What is the EOQ of an item with a project annual demand of 500, and ordering cost of $5, a holding cost percentage of 35%, and an item cost of $10. (Round to the nearest whole unit) For example, 8.52 = 9.
Answer #1
Answer 1
The order point (also called reorder point) is the level of inventory which triggers an action to replenish that particular inventory stock.
It is a minimum amount of an item which a firm holds in stock, such that, when stock falls to this amount, the item must be reordered.
The formula to calculate order point is as under,
Order point = Lead time demand + Safety Stock
Order point = 11 units + 8 units
Order Point = 19 units
Answer 2
Economic order quantity (EOQ) is the order quantity of inventory that minimizes the total cost of inventory management.
The formula to calculate EOQ is as under,
EOQ = SQRT(2 × Annual demand × Cost Per Order / Carrying Cost Per unit)
Annual demand = 500 units
Cost per order = $5
Carrying cost per unit = Unit cost x Holding cost % = $10 x 35% = $3.5 per unit
EOQ = SQRT(2 × 500 × 5 / 3.5)
EOQ = SQRT(1428.57)
EOQ = 37.80
EOQ of an item = 38 Units